A random variable X has the following probability function : x: -2 -1 0 1 2 3 f(x): 0*1 k 0*1 2 2k 0*3 3k (i) Determine the distribution function F(x) of X.

1
Expert's answer
2021-11-15T17:05:12-0500

Solution:

The sum of probabilities in a probability distribution is 1.

So, "0.1+k+0.2+2k+0.3+k=1"

"\Rightarrow 4k+0.6=1
\\\Rightarrow 4k=0.4
\\\Rightarrow k=\dfrac{0.4}4
\\\Rightarrow k=0.1"

Thus, probability distribution is:

"x=-2, f(x)=0.1
\\x=-1, f(x)=0.1
\\x=0,f(x)=0.2
\\x=1,f(x)=0.2
\\x=2,f(x)=0.3
\\x=3,f(x)=0.1"